# MACEval: A Multi-Agent Continual Evaluation Network for Large Models

### v1

@@ Abstract
Hundreds of benchmarks dedicated to evaluating large models from multiple perspectives have been presented over the past few years. Albeit substantial efforts, most of them remain closed-ended and are prone to overfitting due to the potential data contamination in the ever-growing training corpus of large models, thereby undermining the credibility of the evaluation. Moreover, the increasing scale and scope of current benchmarks with transient metrics, as well as the heavily human-dependent curation procedure, pose significant challenges for timely maintenance and adaptation to gauge the advancing capabilities of large models. In this paper, we introduce MACEval, a Multi-Agent Continual Evaluation network for dynamic evaluation of large models, and define a new set of metrics to quantify performance longitudinally and sustainably. MACEval adopts an interactive and autonomous evaluation mode that employs role assignment, in-process data generation, and evaluation routing through a cascaded agent network. Extensive experiments on 9 open-ended tasks with 23 participating large models demonstrate that MACEval is (1) human-free and automatic, mitigating laborious result processing with inter-agent judgment guided; (2) efficient and economical, reducing a considerable amount of data and overhead to obtain similar results compared to related benchmarks; and (3) flexible and scalable, migrating or integrating existing benchmarks via customized evaluation topologies. We hope that MACEval can broaden future directions of large model evaluation.
Figure 1: Paradigm comparison between current large model evaluations and our proposed MACEval.
@@ 1 Introduction
Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) have achieved unprecedented performance in diverse applications such as visual understanding (Achiam et al., 2023; Kavukcuoglu, 2025; Zhu et al., 2025), complex long-text reasoning (math and coding) (Anthropic, 2024; Grattafiori et al., 2024; Team, 2025), and spatial cognition (Wu et al., 2024b; Yang et al., 2024b; Chen et al., 2025d). Evaluation holds an equally pivotal role as pre-training in the development of large models, functioning not only as a benchmark for understanding current performance but also as a critical pillar for aligning model capabilities with intended objectives. In this regard, it is crucial to construct transparent, trustworthy, and zero-contamination evaluations (Zhang et al., 2025a; b).
Despite the proliferation of various evaluation benchmarks, doubts about the genuine capabilities of large models persist (Li et al., 2024d; Yang et al., 2024c), particularly when discrepancies arise between leaderboard scores and real-world user experience. As shown in Fig. 1, we conclude the following three fundamental issues from the perspectives of data source, evaluation pipeline, and evaluation metric: 1) Closed-ended data. Most benchmarks, alongside a corresponding well-organized dataset, which contains evaluation materials (e.g., texts, images, or other modalities) collected from the Internet or existing large-scale datasets, pose potential data contamination risks of overlap with the training set (Xu et al., 2024; Deng et al., 2024). Once constructed, such static datasets are susceptible to obsolescence due to overfitting, especially considering the current rapid monthly update cycle of large models; 2) Human-dependent evaluation pipeline. As the scale and coverage of current benchmarks continue to expand, the drawbacks of heavily human-involved annotation (e.g., query-answer design and data refinement) and result review that traditional benchmarks (Wu et al., 2024a; Liu et al., 2024b; Yue et al., 2024a) rely on are gradually being exposed. Its time-consuming and high-cost characteristics hinder timely maintenance and updates to catch up with the models’ evolving performance; 3) Limited and temporary metric. The current performance quantification strategies that are limited by closed-ended datasets may not reflect the model’s maximum capabilities. Besides, many metrics are transient in nature and fail to incrementally reflect performance changes for future models, making longitudinal assessment challenging. Recently, some researchers have proposed addressing data-related issues by employing techniques such as vision-language bootstrapping (Yang et al., 2024c), establishing specific data generation rules (Zhu et al., 2024a), or meta probing agent (Zhu et al., 2024b). However, solutions for improving evaluation pipelines and metrics remain largely unexplored.
@@
  • We introduce MACEval (Fig. 2), a multi-agent continual evaluation network for dynamic large model assessment. By modeling the evaluation procedure as an autonomous interview, MACEval effectively mitigates data contamination and enhances evaluation efficiency.
  • We propose an AUC-inspired quantitative performance metric for large models to evaluate longitudinally, which can be further incorporated with the evaluation routing to assess the overall performance across various capabilities.
  • We evaluate comprehensively the proprietary and open-source LLMs and MLLMs on several open-ended tasks to demonstrate the effectiveness of the proposed MACEval. Evaluation results show superior efficiency, flexibility, and scalability of MACEval compared to existing benchmarks, thus offering valuable directions for future research on large model evaluation.
@@ 2 Related Work
Benchmarks for Large Models. Benchmarks are the foundations of applied large generative model research. Over the past few years, significant efforts have been made to evaluate LLMs and MLLMs from multiple perspectives (Chang et al., 2024; Li et al., 2024c). From textual comprehension (Bai et al., 2023; Bang et al., 2023) and visual perception (Liu et al., 2023; Fu et al., 2024; Li et al., 2024a; Chen et al., 2025b) to multistep reasoning (Shao et al., 2024; Chen et al., 2024b; Rajabi & Kosecka, 2024) and other specialized domains (Bai et al., 2024; Chen et al., 2025a; Yue et al., 2024a; He et al., 2024; Chen et al., 2025c), evaluation results not only serve as practical guidance for end users but also provide developers with actionable insights for model optimization. With the rapid development of large models and their real-time Internet data crawling strategy used for training, most benchmarks face the risk of being quickly overfitted. Besides, the static, labor-intensive nature and sheer scale of existing benchmarks aggravate the difficulty of maintenance and timely updates (Banerjee et al., 2024), thereby enhancing their susceptibility to data contamination. Therefore, there is an urgent need to establish a long-lasting, dynamic, and robust evaluation framework for large models.
@@
Data Contamination. Recently, data contamination issues have gained widespread attention across evaluation benchmarks for both LLMs and MLLMs (Carlini et al., 2023; Xu et al., 2024; Li & Flanigan, 2024). Researchers from OpenAI and the Llama group conducted contamination studies for GPT-4 (Achiam et al., 2023) and Llama2 (Touvron et al., 2023) on their pre-training data. Study (Li et al., 2024d) reveals significant contamination rates in academic exam-based benchmarks such as MMLU (29.1%) (Hendrycks et al., 2021) and C-Eval (45.8%) (Huang et al., 2023), primarily attributed to the widespread dissemination and circulation of academic test questions. Deng et al. (Deng et al., 2024) proposed a corpora overlap investigation protocol, TS-Guessing, and detected 57% exact match rate of ChatGPT in predicting masked choices in the MMLU test set. Yang et al. (Yang et al., 2024c) reported image overlaps of over 84.4% and 33.2% between SEEDBench (Li et al., 2024a) and pre-training datasets LAION-100M (Schuhmann et al., 2021) and CC3M (Sharma et al., 2018), respectively. Such exposure can significantly impact the reliability of the evaluation, leading to inflated results that do not accurately reflect the true capabilities of the large models. In this paper, we address this problem by reforming the source of evaluation content, introducing AI-generated data and an open-ended task design strategy to enhance benchmark robustness.
Dynamic Evaluation. One recent promising attempt to mitigate the issue of data contamination in large model evaluations is dynamic evaluation. Wang et al. (2025) implemented several perturbations, such as paraphrasing, adding noise, and reversing polarity, to construct evolving instances for testing LLMs against diverse queries and data noise. DyVal (Zhu et al., 2024a) proposes to dynamically generate evaluation samples under pre-defined constraints, which modulates a graph-based algorithm generation structure and fine-grained control over problem difficulty by adjusting the structural complexity. Afterwards, Zhu et al. (2024b) presented meta-probing agents, which automatically refresh an original evaluation problem following psychometric theory on three basic cognitive abilities, including language understanding, problem solving, and domain knowledge. In (Yang et al., 2024c), the authors designed various bootstrapping strategies (e.g. image editing and sentence rephrasing) with complexity control for both image and question modification. However, these studies primarily focus on modifying the sources of evaluation data, without achieving fully dynamic and automatic evaluation processes. Moreover, the corresponding evaluation metrics remain static and transient, failing to adequately reflect the dynamic evolution characteristics of large models.
@@
Figure 2: An overview of our proposed MACEval, which consists of three primary phases: evaluation capability determination, MAEN construction, and open-ended task selection. The pipeline models the evaluation of large models as a multi-round interview process. Specialized agents like interviewers for direct performance evaluation and third-party supervisors for validity assessment of the entire process with a message propagation mechanism, enabling collaboration between interviewer models and other functional models to efficiently and automatically produce reliable evaluation outputs.
@@ 3.1 Design Principles
Focusing on Autonomous Evaluation Process. Unlike existing LLM or MLLM benchmarks that are strongly human-involving in source content collection and relatively independent in cross-ability evaluation, our MACEval follows two basic principles: (1) Not requiring pre-collected evaluation datasets, and all visual or text query-answer pairs are dynamically generated during the process; (2) Progressive capability evaluation scheme with real-time task adjustment. We adhere to the principles in building the MACEval, while making it applicable to evaluating all-round abilities.
@@
Pushing to the Limit. Since existing benchmark datasets are finite and closed-ended, the measured performance scores do not reflect the model’s maximum capabilities. To push the evaluated model to the limit, we adopt a stress-testing strategy in which the model is continuously challenged with increasingly difficult query-answer tasks until it fails to provide a correct response. Henceforth, we can derive a long-standing performance metric by iteratively updating the envelope area formed by connecting performance points across different difficulty levels.
3.2 Curating Open-Ended Tasks
Previous benchmarks for large models primarily focus on collecting raw materials from the Internet or existing general datasets to construct evaluation datasets, leading to potential data contamination risks due to overlaps between evaluation sets and the training data of large models (Touvron et al., 2023; Zhu et al., 2024a; Yang et al., 2024c). In this paper, we avoid such a problem by curating open-ended tasks with quantitatively adjustable difficulty for evaluation and adopt an in-process generation manner through the multi-agent evaluation strategy. This ensures the randomness of the evaluation process while guaranteeing the uniqueness of the target results under the given task. Specifically, we conduct a preliminary exploration of 9 tasks across five key domains currently emphasized in the field: visual perception, textual comprehension, math, algorithms, and coding (White et al., 2025), as listed in Tab. 1. See Appendix B for additional details.
Visual Perception. Evaluating the visual perception capabilities of MLLMs has always been one of the foundational aspects of recent research in MLLMs, featuring prominently in many releases and serving as a key reference for downstream task applications (Bai et al., 2025; Zhu et al., 2025). As illustrated in Fig. 3, we evaluate the low-level and high-level perception capabilities (Wu et al., 2024a) via an Image Quality Perception (IQP) task and a Content Understanding (CU) task, respectively. For the IQP task, we dynamically generate a set of Gaussian white noise with different intensities by increasing the variance and adding it to the original images. For the CU task, we query the MLLM to identify the number of repeated icons in images with progressively increasing grid sizes.
@@ 3.2 Curating Open-Ended Tasks
Figure 3: A data card of 9 open-ended tasks that evaluate the visual perception, text comprehension, math, algorithm reasoning, and coding abilities of large models.
Text Comprehension. For the most fundamental text comprehension task in LLM evaluation, we include a Scrambled Text Understanding (STU) task, i.e., inferring the meaning of a sentence while ignoring typos and misspellings, and a String Parsing (SP) task that evaluates the fine-grained character-level textual processing abilities. Concretely, for the STU task, we disrupt or mask the word structure to generate scrambled text. The proportion of perturbed words relative to the total number of words in the sentence serves as a variable for controlling the difficulty level. For the SP task, the length of the string, namely the number of irrelevant characters, serves as the difficulty variable.
Math. Mathematical reasoning is a cornerstone for assessing the ability of LLMs to resolve complex problems and make multi-hop reasoning, which plays a significant role in general LLM research (Yang et al., 2023; Yue et al., 2024b). Current math-specific benchmarks (Zhang et al., 2024; Xia et al., 2025) mainly collect diverse sets of problems from major textbooks and online resources, such as geometry, linear algebra, and calculus, but rarely provide a quantitative grading of problem difficulty, thereby hindering their direct applicability in the design of open-ended evaluation tasks. Here, we consider the basic arithmetic calculations and encompass two intuitive difficulty control means, i.e., changing the numerical scale and the number of operations, as shown in Fig. 3.
Table 1: Summary of 9 open-ended tasks across 5 capabilities. (IQP: Image Quality Perception; CU: Content Understanding; STU: Scrambled Text Understanding; SP = String Parsing; BTT: Binary Tree Traversal; SPS: Shortest Path Search.)
Capability Task Variable Target
Visual Perception IQP (low-level) Distortion level MLLM
CU (high-level) Grid size MLLM
Text Comprehension STU Disruption ratio LLM
Masking ratio LLM
SP String length LLM
Math Arithmetic Numerical scale LLM
Number of operations LLM
Algorithm BTT Number of nodes (layer) LLM
SPS Number of nodes LLM
Edge density LLM
Coding Generation Cyclomatic complexity LLM
Debugging Missing ratio LLM
Algorithm. As a derivative task of basic mathematics, solving algorithmic problems constitutes another important aspect for assessing the reasoning capabilities of LLMs. Specifically, we select two common tasks in computer science, Binary Tree Traversal (BTT) and Shortest Path Search (SPS), for evaluation. Among them, the number of nodes and the edge density that determine the complexity of the data structure are taken as the variables.
Coding. Coding is among the most practical capabilities of LLMs in real-world scenarios. Following (Chen et al., 2024c; Jain et al., 2025), we design two open-ended code generation and debugging tasks with dynamic difficulty. The generation task specifically focuses on the ability of models to parse a cyclomatic complexity evolved coding question statement and write an executable answer. The debugging scenario assesses the self-repair capabilities of LLMs. Here, the model is given a problem statement and its corresponding code with partially removed content (missing ratio), and then prompted to generate a repaired version.
@@ 3.3 Definition and Formulation
We define the input space of a multi-agent system as $\mathbb{I}=(\mathbb{M},\mathbb{R},\mathbb{T})$, where $\mathbb{M}$ denotes the candidate pool of LLMs or MLLMs. $\mathbb{R}$ represents the set of pre-defined agent roles (e.g., interviewee, interviewer, and supervisor), and $\mathbb{T}$ is the set of evaluation topologies, i.e. collaboration modes, such as point-to-point link, tree, star, and multi-hop network. Within the evaluation space, a Multi-Agent Evaluation Network (MAEN) instance is defined as follows:
Definition 1 (Multi-Agent Evaluation Network) The graph-based MAEN $\mathcal{G}$ includes several large models with distinct identities (node type), participating collaboratively (edge type) in a configurable (topological type) evaluation process:
$\mathcal{G}=\left\{{\{{\mathcal{M}_{i}}\}_{i=1}^{M},\{{\mathcal{R}_{j}}\}_{j=1}^{R},\{{\mathcal{T}_{k}}\}_{k=1}^{T}}\right\},\;\;{\mathcal{M}_{i}}\in\mathbb{M},\;{\mathcal{R}_{j}}\in\mathbb{R},\;{\mathcal{T}_{k}}\in\mathbb{T},$ (1)
where $\mathcal{M}$, $\mathcal{R}$, and $\mathcal{T}$ correspond to the selected large model, role function, and evaluation network structure, respectively. $M$, $R$, and $T$ are the number of each corresponding component.
Definition 2 (Multi-Agent Continual Evaluation Stream) The MACEval stream can be denoted by a mapping function $f:\mathbb{M}\times\mathbb{R}\times\mathbb{T}\to\mathcal{G}$ that maps the input space $\mathbb{I}$ to an MAEN $\mathcal{G}$ tailored for the evaluated capability $\mathcal{C}$ from a non-stationary data stream $\mathcal{S}=\left\{{{{({q,a})}_{0}},{{({q,a})}_{1}},\cdots,{{({q,a})}_{\infty}}}\right\}$, where data tuple $({q,a})_{t}$} (query-answer pair) is dynamically generated via a step $t$-dependent function.
@@
Specifically, when the message passing terminates at a supervisor node, the aggregated messages store the information of the entire MAEN, which reflects the validity of the evaluation process.
3.4 A Framework for MACEval

### v2

@@ Abstract
Hundreds of benchmarks dedicated to evaluating large models have been presented over the past few years. However, most of them remain closed-ended and are prone to overfitting due to the potential data contamination. Moreover, the increasing scale and scope of current benchmarks with transient metrics, as well as the heavily human-dependent curation procedure, pose significant challenges for timely maintenance and adaptation. In this paper, we introduce MACEval, a Multi-Agent Continual Evaluation network for dynamic evaluation of large models, and define new metrics to quantify performance longitudinally. MACEval employs an interactive and autonomous evaluation mode, utilizing role assignment, in-process data generation, and evaluation routing through a cascaded agent network. Extensive experiments on 23 large models demonstrate the effectiveness of MACEval, which also lightens the evaluation process and reduces a considerable amount of overhead. We hope that MACEval can broaden future directions of large model evaluation. Project page: https://github.com/zijianchen98/MACEval.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) have achieved unprecedented performance in diverse applications such as visual understanding (Achiam et al., 2023; Kavukcuoglu, 2025; Zhu et al., 2025; Chen et al., 2025), complex long-text reasoning (math and coding) (Anthropic, 2024; Grattafiori et al., 2024; Team, 2025), and spatial cognition (Wu et al., 2024b; Yang et al., 2024b). Evaluation holds an equally pivotal role as pre-training in the development of large models, functioning not only as a benchmark for understanding current performance but also as a critical pillar for aligning model capabilities with intended objectives. In this regard, it is crucial to construct transparent, trustworthy, and zero-contamination evaluation processes.
Figure 1: Paradigm comparison between current large model benchmarks and the proposed MACEval.
Despite the proliferation of various evaluation benchmarks, doubts about the genuine capabilities of large models persist (Li et al., 2024c; Yang et al., 2024c), particularly when discrepancies arise between leaderboard scores and real-world user experience. As shown in Fig. 1, we conclude the following three fundamental issues from the perspectives of data source, evaluation pipeline, and evaluation metric: 1) Closed-ended data. Most benchmarks, alongside a corresponding well-organized dataset, which contains evaluation materials (e.g., texts, images, or other modalities) collected from the Internet or existing large-scale datasets, pose potential data contamination risks of overlap with the training set (Xu et al., 2024; Deng et al., 2024). Once constructed, such static datasets are susceptible to obsolescence due to overfitting, especially considering the current rapid monthly update cycle of large models. 2) Human-dependent evaluation pipeline. As the scale and coverage of current benchmarks continue to expand, the drawbacks of heavily human-involved annotation (e.g., query-answer design and data refinement) and result review that traditional benchmarks (Wu et al., 2024a; Liu et al., 2024b; Yue et al., 2024) rely on are gradually being exposed. Its time-consuming and high-cost characteristics hinder timely maintenance and updates to catch up with the models’ evolving performance. 3) Limited and temporary metric. The current performance quantification strategies that are limited by closed-ended datasets may not reflect the model’s maximum capabilities. Besides, many metrics are transient in nature and fail to incrementally reflect performance changes for future models, making longitudinal assessment challenging. Recently, some researchers have proposed to address data-related issues by employing techniques such as vision-language bootstrapping (Yang et al., 2024c), establishing specific data generation rules (Zhu et al., 2024a), or meta probing agent (Zhu et al., 2024b). However, solutions for improving evaluation pipelines and metrics remain largely unexplored.
@@
  • We introduce MACEval, a multi-agent continual evaluation network for dynamic large model assessment. By modeling the evaluation procedure as an autonomous interview, MACEval effectively mitigates data contamination and enhances evaluation efficiency.
  • We propose an AUC-inspired metric for large models to evaluate longitudinally, which can be incorporated with the evaluation routing to assess the overall performance utilization rate.
  • We evaluate the proprietary and open-source LLMs and MLLMs on several open-ended tasks to demonstrate the effectiveness of the proposed MACEval. Evaluation results demonstrate the superior efficiency, flexibility, and scalability of MACEval compared to existing benchmarks, providing valuable insights for future research on large model evaluation.
@@ 2 Related Work
Data Contamination. Recently, data contamination issues have gained widespread attention across evaluation benchmarks for both LLMs and MLLMs (Carlini et al., 2023; Xu et al., 2024; Li & Flanigan, 2024). Researchers from OpenAI and the Llama group conducted contamination studies for GPT-4 (Achiam et al., 2023) and Llama2 (Touvron et al., 2023) on their pre-training data. Study (Li et al., 2024c) reveals significant contamination rates in academic exam-based benchmarks such as MMLU (29.1%) (Hendrycks et al., 2021) and C-Eval (45.8%) (Huang et al., 2023), primarily attributed to the widespread dissemination and circulation of academic test questions. Deng et al. (Deng et al., 2024) proposed a corpora overlap investigation protocol, TS-Guessing, and detected 57% exact match rate of ChatGPT in predicting masked choices in the MMLU test set. Yang et al. (Yang et al., 2024c) reported image overlaps of over 84.4% and 33.2% between SEEDBench (Li et al., 2024a) and pre-training datasets LAION-100M (Schuhmann et al., 2021) and CC3M (Sharma et al., 2018), respectively. Such exposure can significantly impact the reliability of the evaluation, leading to inflated results that do not accurately reflect the true capabilities of the large models. In this paper, we address this problem by reforming the source of evaluation content, introducing AI-generated data, and an open-ended task design strategy to enhance benchmark robustness.
@@
Figure 2: An overview of our proposed MACEval, which consists of three primary phases: evaluation capability determination, MAEN construction, and open-ended task selection. The pipeline models the evaluation of large models as a multi-round interview process.
Dynamic Evaluation. One recent promising attempt to mitigate the issue of data contamination in large model evaluations is dynamic evaluation. (Wang et al., 2025) implemented several perturbations, such as paraphrasing, adding noise, and reversing polarity, to construct evolving instances for testing LLMs against diverse queries and data noise. DyVal (Zhu et al., 2024a) proposes to dynamically generate evaluation samples under pre-defined constraints, which modulates a graph-based algorithm generation structure and fine-grained control over problem difficulty by adjusting the structural complexity. Afterwards, (Zhu et al., 2024b) presented meta-probing agents, which automatically refresh an original evaluation problem following psychometric theory on three basic cognitive abilities, including language understanding, problem solving, and domain knowledge. In (Yang et al., 2024c), the authors designed various bootstrapping strategies (e.g., image editing and sentence rephrasing) with complexity control for both image and question modification. However, these studies primarily focus on modifying the sources of evaluation data, without achieving fully dynamic and automatic evaluation processes. Moreover, the corresponding evaluation metrics remain static and transient, failing to reflect the dynamic evolution characteristics of large models adequately.
@@ 3.1 Design Principles
Focusing on Autonomous Evaluation Process. Unlike existing LLM or MLLM benchmarks that are strongly human-involving in source content collection and relatively independent in cross-ability evaluation, our MACEval follows two basic principles: (1) Mitigating the need for pre-curated evaluation sets, all visual and textual query–answer pairs are generated on-the-fly; We conduct a preliminary exploration on 9 tasks across five key domains currently emphasized in the field, i.e., visual perception, textual comprehension, math, algorithms, and coding (White et al., 2025; Zhang et al., 2025). As illustrated in Fig. 3, each task incorporates dynamic difficulty variables (e.g., distortion level, string length, or cyclomatic complexity) to enable quantitative assessment. (2) Progressive capability evaluation scheme with real-time task adjustment. We adhere to the principles in building the MACEval, while making it applicable to evaluating all-round abilities. More details are in App. A.
@@
Pushing to the Limit. Since existing benchmark datasets are finite and closed-ended, the measured performance may not reflect the model’s maximum capabilities. To push the evaluated model to the limit, we adopt a stress-testing strategy, where the model is continuously challenged with increasingly difficult query-answer tasks until it fails to provide a correct response. Henceforth, we can derive a long-standing metric by iteratively updating the envelope area formed by connecting performance points across different difficulty levels.
@@
Figure 3: A data card of 10 open-ended tasks that evaluate the visual perception, text comprehension, math, algorithm reasoning, and coding abilities of large models. The difficulty variable of each task is indicated in red.
@@ 3.2 Definition and Formulation
We define the input space of a multi-agent system as $\mathbb{I}=(\mathbb{M},\mathbb{R},\mathbb{T})$, where $\mathbb{M}$ denotes the candidate pool of LLMs or MLLMs. $\mathbb{R}$ denotes the set of pre-defined agent roles (e.g., interviewee, interviewer, and supervisor), and $\mathbb{T}$ is the set of evaluation topologies, i.e. collaboration modes, such as point-to-point link, tree, star, and multi-hop network. Within the evaluation space, a Multi-Agent Evaluation Network (MAEN) is defined as follows:
Definition 1 (Multi-Agent Evaluation Network): The MAEN $\mathcal{G}$ includes several large models with distinct identities (node type $\mathcal{M}$), participating collaboratively (edge type $\mathcal{R}$) in a configurable (topological type $\mathcal{T}$) evaluation process:
$\mathcal{G}=\left\{{\{{\mathcal{M}_{i}}\}_{i=1}^{M},\{{\mathcal{R}_{j}}\}_{j=1}^{R},\{{\mathcal{T}_{k}}\}_{k=1}^{T}}\right\},$ (1)
where $M$, $R$, and $T$ are the number of each corresponding component.
Definition 2 (Multi-Agent Continual Evaluation Stream): The MACEval stream can be denoted by a mapping function $f:\mathbb{M}\times\mathbb{R}\times\mathbb{T}\to\mathcal{G}$ that maps the input space $\mathbb{I}$ to an MAEN $\mathcal{G}$ tailored for the evaluated capability $\mathcal{C}$ from a non-stationary data stream $\mathcal{S}=\left\{{{{({q,a})}_{0}},{{({q,a})}_{1}},\cdots,{{({q,a})}_{\infty}}}\right\}$, where data tuple $({q,a})_{t}$} (query-answer pair) is dynamically generated via a step $t$-dependent function.
@@
Specifically, when the message passing terminates at a supervisor node, the aggregated messages store the information of the entire MAEN.
3.3 MACEval Framework